Bent Gregorian Technical Observing : 08 September 2009 UT

Observer: J. Hill (mtn), D. Miller (Tucson)
Telescope Operator: W. Wack
SW support: N. Cushing

Summary

We opened at sunset and worked through mostly clear sky until 2AM. Seeing was excellent between 0.5 and 0.8 arcsec.

The first significant problem is that while flipped PCS/GCS (as LUCIFER) can find the guide star off-axis, the AGW coordinate transformation is screwed up at the 5 arcsec level by the flip. We took lots of data on StoneMagw to demonstrate that. Jesper told us what to flip in oacontrol.conf, but we weren't able to test that on sky.

The next significant problem is that unflipped PCS/GCS (as IRTC) cannot locate guide stars which are off the X-axis. Thus, it seems we have a new X-flip (compared to B29) that nobody will admit to. Guiding is OK, so the cure has to move the AGW probe without changing guiding. See email from Michele summarizing how the various flips and transformations differ.

The third significant problem is that the PCS/GCS X-flip for LUCIFER messes up the pupil shift calculation in GCS, so we can do WFS off the Y-axis. GCS has a parameter to change this, but it seems not to have any effect.

The next significant problem was that Michele copying logfiles on lbtmu304 could slow PCS to the point of polynomials arriving 4 seconds late. This stops tracking after they are 1 second late.

Doug spent alot of time updating the IDL code to match the various and sundry TCS changes such as new DD variable names.

We spent an hour of cloudy time learning that you have to wait 5-6 minutes after starting oacserver before you can startAGW. We need to shorten the timeouts for AGWs not on the mountain.

Preparations

Restarted oacserver and chk4nport this afternoon to untangle some RPC problem.

Fixed the call for getdat in home$transtar.cl

Took the fogged plexiglass window out of the axis200p webcam enclosure, and replaced it with kitchen wrap. This made a big improvement in the image. Was unable to get the replacement camera running.

Details

Open

01:45 Open shutter doors at sunset. Yes, really at sunset. Sky has a fair bit of scattered cloud, but its better than we have seen in a while. T=10 degC, D=6 degC. Wind calm.

Home the AGw probe motors.

02:01 pointto WT10_268 TelM="ACQUIRE" PAM- gs=0 Couldn't set full frame ROI. Continuing in the hope the image is large enough anyway.
Called Norm.

02:16 pointto WT10_268 TelM="ACQUIRE" PAM- gs=0 - same error.

02:17 Stop and restart GCS

02:18 pointto WT10_268 TelM="ACQUIRE" PAM- gs=0 - no AGW unit selected. (This is still a bug in GCS.)

As LUCIFER

02:19 authorize LUCIFER again.

02:20 pointto WT10_268 TelM="ACQUIRE" PAM- gs=0 - Adjust IE=-87 CA=-39 (so restarting GCS did the trick.)

02:27 pointto BS9177 TelM="ACQUIRE" PAM- - Adjust IE=-92 CA=-34 (on the other side of the sky)

02:34 pointto BS9177 ACTIVE PAM- - After a bit of active optics, seeing as measured on guider image is 0.5 arcsec!

02:40 pointto BS9177 ACTIVE gs=4 - couldn't move probe there.

02:40 pointto BS9177 ACTIVE gs=3 (can't guess a good guide star without a finder chart)

02:45 pointto BS9183 ACTIVE gs=1 - couldn't move probe there.

02:48 pointto BS9183 ACTIVE gs=0 - no star in acquisition image

02:50 pointto M5_1526 TelM="ACQUIRE" PAM- star is off the left edge - Adjust IE=-92 CA=-58

02:54 pointto StoneMagw ACTIVE gs=55 - star was at very edge of acquisition image at ~2:00. Was it the correct star?

02:58 pointto StoneMagw TelM="ACQUIRE" gs=0 PAM- - Adjust IE=-99 CA=-70

03:02 pointto StoneMagw ACTIVE gs=0 PA=0 - OK

03:04 pointto StoneMagw ACTIVE gs=0 PA=90 - OK

03:05 pointto StoneMagw TelM="ACTIVE" gs=51 PA=0 - missed by ~4 arcsec.

03:08 pointto StoneMagw TelM="ACTIVE" gs=45 PA=180 - missed by ~3 arcsec.

03:12 task transtar=home$transtar.cl

03:13 cl < tran_StoneM_00.cl with 1 sec x 40 IRTC cubes (1 sec was too long)

transtar (39) (missed the IRTC image because I was in video mode)
transtar (40) irtc.20090906_00004_cube.fit (note wrong date in the name)
transtar (51)
transtar (72)
transtar (68)
transtar (58) irtc.20090908_00008_cube.fit
transtar (52)
transtar (48)
transtar (44)
transtar (43)
transtar (38)
transtar (37)
transtar (35)
transtar (26)
transtar (20) irtc.20090908_00017_cube.fit
transtar (13) (Chase failed to make tolerance - image was not guided.)
transtar (18)
transtar (51)

Conclude that the transform is screwed up. Central star is moving around on IRTC by +-5 arcsec.

As IRTC

03:52 Authorize as IRTC, register IRTC

03:53 pointto StoneMagw TelM="ACQUIRE" gs=0 PAM- - Adjust IE=-77 CA=-46

03:57 pointto StoneMagw TelM="ACTIVE" gs=51 PA=0 - started guiding on some very faint star. We must have centered on the wrong one.

04:00 pointto M5_0746 TelM="ACQUIRE" PAM- gs=0 Adjust IE=-73 CA=-38

04:05 pointto StoneMagw TelM="ACTIVE" gs=51 PA=0 - no star found - could it be clouds?

04:08 pointto StoneMagw TelM="ACTIVE" gs=0 PA=0 - off by 10 arcsec in the acquisition image ??

04:09 Clear secondary offload, and make a bootstrap convergence test. WFS converges quite well and removes ~6000 nm of aberration in 4 iterations.

04:12 pointto StoneMagw TelM="ACQUIRE" PAM- gs=0 - Adjust IE=-77 CA=-48

04:16 pointto StoneMagw TelM="ACTIVE" gs=51 PA=0 - we don't find the star so, I'm confused.

Called Michele.

04:32 Made IRTC flips false false rather than false true that they were set on Monday. Restart PCS.

04:33 pointto StoneMagw TelM="ACQUIRE" PAM- gs=0 - star had drifted 5 arcsec in 20 minutes open-loop. Adjust IE=-80 CA=-51

04:36 pointto StoneMagw TelM="ACTIVE" gs=0 PA=0 - guiding runs away -

04:38 Put the flips back and restart PCS.

04:39 pointto StoneMagw TelM="ACTIVE" gs=0 PA=0

04:39 pointto StoneMagw TelM="ACTIVE" gs=53 PA=INDEF=99.64 - started guiding, but then failed with no guide star detected warning ??

04:43 pointto StoneMagw TelM="ACTIVE" gs=53 - failed for bad poly on rotator ?? What happened here? Lost 6 minutes.

04:49 pointto StoneMagw TelM="ACTIVE" gs=53 - OK for a while, but then bad polys again on all axes. Was it because Michele was making extra extra activity on lbtmu302?

04:53 pointto StoneMagw TelM="ACTIVE" gs=55 PA=180 - no bright star in acquisition image.

04:56 pointto StoneMagw TelM="ACTIVE" gs=0 PA=0 - Yes, stale polys appear when Michele copies a big log file.

05:02 pointto StoneMagw TelM="ACTIVE" gs=55 PA=INDEF=142.94 - OK - Images are 0.7 to 0.8 arcsec FWHM at EL=35.

05:05 pointto StoneMagw TelM="ACTIVE" gs=55 PA=162 forcing the probe off the Y-axis. - no star in acquisition image.

As Lucifer

05:08 Authorize as LUCIFER

05:08 Register as LUCIFER, handing driving over to Doug.

05:15 Send pointto BS9188 TelMode=ACQUIRE gs=0

05:16 Wayne cleared M2 corrections: Focus and coma appear (John comments that the collimation model was better for focus at the beginning of the night

05:18 Wayne set IE=-101 CA=-46

05:20 John add focus of -0.4 mm Zglobal on primary to get better image

05:20 Humidity spike from 80 to 88. Cloud north of Safford

05:25 Send pointto BS9188 TelMode=ACTIVE gs=0

05:45 GCS collimate then pointto BS9188 TelMode=GUIDE gs=7 PA=276

06:00 There is not an LBTO account on the GCS machine, lbtmu304. Will set up telescope account to run the IDL routines (and make an IssueTrak). John is taking over. Confirming again that off-axis stars are correct for LUCIFER.

06:00 pointto StoneOagw ACQUIRE gs=0 PAM- Adjust IE=-100 CA=-54

Seeing on guider is 1 arcsec

06:10 pointto StoneOagw TelM="ACTIVE" gs=1 PA=INDEF=299.24 - star is 5 arcsec to right on acq image. Guiding OK.

06:13 pointto StoneOagw TelM="ACTIVE" gs=1 PA=309.24 - OK

06:15 pointto StoneOagw TelM="ACTIVE" gs=1 PA=329.24 - OK

06:16 pointto StoneOagw TelM="ACTIVE" gs=1 PA=349.24 - OK

offset 5 0 ABS RADEC

offset 0 5

offset 0 0

Wavefront sensing is goofy here. Seems to be diverging. Probably the pupil shift has to flip X also.

06:24 pointto StoneOagw TelM="ACTIVE" gs=1 PA=INDEF=299.24 - That allows it to recover.

06:28 change WFS_pupil_shift_x from +1.0 to -1.0 in AIP_L.cfg

06:30 pointto StoneOagw TelM="ACTIVE" gs=1 PA=349.24 - but WFS still diverges.

06:34 pointto StoneOagw TelM="ACTIVE" gs=1 PA=INDEF=299.24

06:39 stop guiding, stop and restart GCS

06:40 pointto StoneOagw TelM="ACTIVE" gs=1 PA=349.24 - no AGW unit selected

06:41 selectAGW AIP_L.cfg

06:42 pointto StoneOagw TelM="ACTIVE" gs=1 PA=349.24 - still diverging, the parameter for flipping the pupil shift didn't take.

06:44 pointto StoneOagw TelM="ACTIVE" gs=1 PA=INDEF=299.24

06:50 Doug takes over: Field data for Andrew. Now have the IDL routines running on the telescope account on lbtmu304. Preset to ZBS9188. Wayne set IE=-101 CA=-66

07:05 preset in Active mode on gs=0. Not convinced that we have the correct star. Will collimate and then guide offaxis to see if we find the off-axis star.

07:08 preset in Active mode on gs=7. Star not in acquisition image

07:15 preset to Stone0 fields in ACQUIRE mode

07:20 Wayne put Stone0 star on hotspot IE=-108 CA=-62

07:23 preset ACTIVE to StoneO, gs=7 (PA=269)

07:35 tried to run agw_collimate. Permission problems in /Repository (fixed). Old version of IDL routines could not read DD variables from IIF (fixed, updated idl/wfsc routines)

07:45 tried agw_collimate again. Troubles reading DD variables with IDL. John takes over while Doug finds the problem.

Stone0agw
GS PA WFSC
0 260.62  
7 269.62  
0 260.62  
     

Close

08:10 Close for clouds

08:21 Closed dome preset to StoneAagw using local catalog for LBTO home$LBTtools/IRTC/Catalogs/LBTdefault.cat

08:30 changed pcsmx from -1 to +1 in /etc/oacontrol/oacontrol.conf to flip the transform.

08:32 stopAGW -u 3

08:34 stop and restart oacserver. stop and restart GCS. startAGW -u 3 (RPC error)

08:37 stop all of oacserver, chk4nports, npreal2d to cure RPC error.

08:45 stop all and try again

08:52 sudo reboot lbtmu105, but that didn't help.

The problem was that we weren't waiting long enough. It took nearly 6 minutes for the oacserver to start. Each non-existing AGW seems to have a 3-minute timeout.
Sep  8 09:09:52 lbtmu105 oacserver[8432]: LOG_INFO starting ...
Sep  8 09:13:01 lbtmu105 oacserver[8432]: LOG_ERR Could not connect to PMAC #0 
Sep  8 09:16:10 lbtmu105 oacserver[8432]: LOG_ERR Could not connect to PMAC #1 
Sep  8 09:16:13 lbtmu105 oacserver[8432]: LOG_ERR Could not connect to PMAC #2 
Sep  8 09:16:16 lbtmu105 oacserver[8432]: LOG_ERR Could not connect to PMAC #3 
Sep  8 09:16:17 lbtmu105 oacserver[8432]: LOG_ERR getMotorErrors(): Error while flushing PMAC 
Sep  8 09:16:17 lbtmu105 oacserver[8432]: LOG_ERR getMotorErrors(): pthread_exit() 
Sep  8 09:16:17 lbtmu105 oacserver[8432]: LOG_ERR geterrors WD [5]: Thread exited [agw #0] 
Sep  8 09:16:17 lbtmu105 oacserver[8432]: LOG_ERR getMotorErrors(): Error while flushing PMAC 
Sep  8 09:16:17 lbtmu105 oacserver[8432]: LOG_ERR getMotorErrors(): pthread_exit() 
Sep  8 09:16:17 lbtmu105 oacserver[8432]: LOG_ERR Fatal following error exceeded WD [26]: Thread exited [agw #1] 

09:32 startAGW -u 3 - succeeded

09:34 restart GCS

09:43 pointto with closed dome.

-- JohnHill - 08 Sep 2009
Topic revision: r7 - 09 Sep 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback